# Noise Synthesis for Low-Light Image Denoising with Diffusion Models

### v1

@@ C.1 Network Architecture
Positional Encoding.
To integrate pixel coordinate information into our diffusion network, we propose to use positional encoding. We first use sinusoidal positional encoding [47] to encode the pixel coordinates as
@@
Camera Setting Encoding.
Each camera setting’s corresponding embedding $\mathbf{z}$ is integrated with the image feature $\mathbf{f}$ using cross-attention as
@@ C.3 Training Configurations
Diffusion Model Training.
Before training the diffusion model, we perform data resampling: for camera settings with fewer than 100 training pairs, we augment the data by repeating samples until the count reaches 100. We train our diffusion model using the Adam optimizer for 1,000 epochs. The learning rate is initially set to $41e-$ 4 and decays to $51e-$ 5 by the end of the training, using a cosine annealing schedule. The batch size is set to 12. We randomly crop $512\times 512$ patches from the training samples. All Sony training pairs from the SID dataset are used for training, with their ISO values and exposure time ratios encoded.
@@
Denoising Network Training.
To assess the quality of the synthetic data, we train denoising networks using the $L1$ loss between the network’s output and the reference clean image, then evaluate their denoising performance. For a fair comparison, we use the same U-Net architecture employed in [11]. The denoising networks are trained using the Adam optimizer for 500 epochs. During training, the batch size is 4, and we randomly crop $256\times 256$ patches, applying random horizontal flips for data augmentation.
C.4 Clean-Noisy Pair Generation
@@
Appendix D Evaluation Protocol: Noise Mean and Variance
@@
Figure S2: Ablation study on influence of training-data diversity for denoising. Paired real data: The denoising network is trained using real data from the SID Sony training set [11], where some camera settings have very few unique scenes (e.g., the camera setting with ISO 6400 and ratio 300 has only one unique scenes). Ours: The denoising network is trained with our full set of synthetic data, where camera settings with limited unique scenes are augmented by scenes from other settings. Synthetic data is generated using the augmented scenes. Ours (Limited data): The denoising network is trained with limited synthetic data, where the number of scenes for each camera setting matches the original SID training set. The above three test samples from the SID testing set: (1) ISO 6400, ratio 300 (only one unique clean image in the original SID training set); (2) ISO 12800, ratio 100 (three unique clean images originally); and (3) ISO 800, ratio 100 (two unique clean images originally). It can be observed that the denoising performance improves with more synthetic data.
@@
Appendix E More Discussions
E.1 Influence of Training-Data Diversity
@@
Standard Configuration.
The original SID training set contains varying numbers of scenes across different camera settings, with some settings having very few scenes. For instance, the camera setting with ISO 6400 and ratio 300 has only one unique scenes. With our data generation model, however, we can leverage all the clean images in the SID training set to generate numerous training pairs for each camera setting. Specifically, if there are $K$ clean images in total, we can generate $K$ noisy images for each camera setting by conditioning our diffusion model on the camera setting embedding. Furthermore, since DDPM sampling is non-deterministic, different random seeds allow us to generate infinite noisy images from a single clean image. For efficiency, though, we generate a certain number of noisy images for each camera setting. For any setting with fewer than 6 unique clean images ($K_{i}<6$), we randomly sample $30-K_{i}$ additional clean images from other settings to augment the clean image number, then generate noisy images from both the original and augmented clean images. For settings with more than 6 clean images, we use only the original clean images for data generation.
Limited-Data Configuration.
Here, we generate data without clean image augmentation. Specifically, for each camera setting with $K_{i}$ clean images in the SID training set, we generate noisy images using only these $K_{i}$ clean images. Therefore, the denoising network under this limited-data configuration uses the same number of scenes per camera setting as in the original SID training set.
@@
Fig. S2 presents a visual comparison using three test samples from the SID testing set: (1) ISO 6400, ratio 300 (only one unique clean image in the original SID training set); (2) ISO 12800, ratio 100 (three unique clean images originally); and (3) ISO 800, ratio 100 (two unique clean images originally). As shown, Ours produces better results than Ours (Limited data), benefiting from the additional synthetic training data. This further demonstrates that denoising performance improves with more synthetic data.
@@ E.2 Sampling Method
E.3 Noise Decomposition
Implicit Noise Decomposition.
While noise decomposition on real noise images is not possible due to the absence of ground-truth for the signal-dependent and signal-independent components, the conditional generation capability of our diffusion model allows us to perform implicit noise decomposition on the generated noise.
@@ E.3 Noise Decomposition
We further analyze the statistics of different noise components. In Fig. S4 (g) and (j), we compare the variance and mean values of our generated noise with those of the real noise; they are calculated using the method described in Sec. D. The results indicate that our method produces a noise distribution similar to the real noise. In Fig. S4 (h) and (k), we present the statistics of the signal-independent noise; they clearly show that the noise mean and variance are independent of the clean pixel values. Conversely, Fig. S4 (i) and (l) depict the statistics of the signal-dependent noise, demonstrating that both the noise variance and mean are correlated with the clean pixel values.
@@
Additional Evidence Supporting the Effectiveness of the Two-Branch Network.
Meanwhile, we also apply the same decomposition process to the noise generated by the UNet-only architecture, with the results shown in Fig.S4 (m)-(u). It is evident that the signal-independent and signal-dependent components are not fully disentangled. For instance, in Fig.S4 (o), low-frequency noise is still present at the bottom of the signal-dependent component. Moreover, the we can clear This performance difference between the two architectures further highlights the effectiveness of the two-branch architecture in modeling signal-independent and signal-dependent noise, simplifying the network’s learning process.
Figure S6: Histogram of noise values and Kullback-Leibler (KL) divergence between generated noise and real noise for four ablation variants: the UNet-only architecture, model without positional encoding (W/o PE), our full model with DDIM sampling (Ours (DDIM)) and our full model with DDPM sampling (Ours). Evaluations are conducted across three different ISO settings. The results produce by our full model are closest to the real noise distribution.
E.4 Noise Histogram
@@
Appendix F More Visual Comparisons
@@
In this section, we provide more visual comparisons of denoising results from different networks trained on data synthesized using different methods. Fig. S7 - Fig. S11 are examples from the SID test set and Fig. S12 - Fig. S13 are examples from the ELD test set. Our method demonstrates effectiveness in handling row noise, as evidenced in the second and third examples of Fig.S7 and the first example of Fig.S8. Additionally, it excels at mitigating low-frequency noise, particularly visible in the second and third examples of Fig.S9. Moreover, our approach enhances detail retention in denoising results, for example, producing clearer text compared to other methods, as shown in the second and third examples of Fig.S9 and throughout Fig. S12.
Appendix G Ethics Considerations
This paper addresses the task of noise synthesis for image denoising, primarily aimed at enhancing image quality in low-light photography. While the technique can improve photographic quality for general use, it also has implications for fields such as surveillance and medical imaging, where denoising and noise synthesis could potentially impact data integrity if misused. We recognize the importance of transparency in synthetic data applications and encourage responsible use to avoid unintended consequences, such as creating misleadingly “cleaned” images in sensitive areas where accurate data representation is essential.
No known negative societal impacts are associated with this research, and we encourage users to apply these methods ethically.
@@ Appendix G Ethics Considerations
Figure S7: Visual comparisons of denoising results from networks trained on data synthesized using different methods, alongside comparisons with a model trained on real clean-noisy pairs from the SID training set. All test samples are from the SID Sony test set. Results here are processed by the same gamma correction ($\textbf{I}_{out}=\textbf{I}_{in}^{1/1.6}$) for a clearer view as the main paper.
@@
Figure S8: Visual comparisons of denoising results from networks trained on data synthesized using different methods, alongside comparisons with a model trained on real clean-noisy pairs from the SID training set. All test samples are from the SID Sony test set. Results here are processed by the same gamma correction ($\textbf{I}_{out}=\textbf{I}_{in}^{1/1.6}$) for a clearer view as the main paper.
@@
Figure S9: Visual comparisons of denoising results from networks trained on data synthesized using different methods, alongside comparisons with a model trained on real clean-noisy pairs from the SID training set. All test samples are from the SID Sony test set. Results here are processed by the same gamma correction ($\textbf{I}_{out}=\textbf{I}_{in}^{1/1.6}$) for a clearer view as the main paper.
@@
Figure S10: Visual comparisons of denoising results from networks trained on data synthesized using different methods, alongside comparisons with a model trained on real clean-noisy pairs from the SID training set. All test samples are from the SID Sony test set. Results here are processed by the same gamma correction ($\textbf{I}_{out}=\textbf{I}_{in}^{1/1.6}$) for a clearer view as the main paper.
Figure S11: Visual comparisons of denoising results from networks trained on data synthesized using different methods, alongside comparisons with a model trained on real clean-noisy pairs from the SID training set. All test samples are from the SID Sony test set. Results here are processed by the same gamma correction ($\textbf{I}_{out}=\textbf{I}_{in}^{1/1.6}$) for a clearer view as the main paper.
@@
Figure S12: Visual comparisons of denoising results from networks trained on data synthesized using different methods, alongside comparisons with a model trained on real clean-noisy pairs from the SID training set. All test samples are from the ELD test set. Results here are processed by the same gamma correction ($\textbf{I}_{out}=\textbf{I}_{in}^{1/1.6}$) for a clearer view as the main paper.
@@
Figure S13: Visual comparisons of denoising results from networks trained on data synthesized using different methods, alongside comparisons with a model trained on real clean-noisy pairs from the SID training set. All test samples are from the ELD test set. Results here are processed by the same gamma correction ($\textbf{I}_{out}=\textbf{I}_{in}^{1/1.6}$) for a clearer view as the main paper.

### v2

@@ 9.1 Network Architecture
Positional Encoding. To integrate pixel coordinate information into our diffusion network, we propose to use positional encoding. We first use sinusoidal positional encoding [55] to encode the pixel coordinates as
@@
Camera Setting Encoding. Each camera setting’s corresponding embedding $\mathbf{z}$ is integrated with the image feature $\mathbf{f}$ using cross-attention as
@@ 9.3 Training Configurations
Diffusion Model Training. Before training the diffusion model, we perform data resampling: for camera settings with fewer than 100 training pairs, we augment the data by repeating samples until the count reaches 100. We train our diffusion model using the Adam optimizer for 1,000 epochs. The learning rate is initially set to $41e-$ 4 and decays to $51e-$ 5 by the end of the training, using a cosine annealing schedule. The batch size is set to 12. We randomly crop $512\times 512$ patches from the training samples. All Sony training pairs from the SID dataset are used for training, with their ISO values and exposure time ratios encoded.
@@
Denoising Network Training. The denoising networks are trained using the Adam optimizer for 500 epochs. During training, the batch size is 4, and we randomly crop $256\times 256$ patches, applying random horizontal flips for data augmentation.
9.4 Clean-Noisy Pair Generation
@@
10 Evaluation Protocol: Noise Mean and Variance
@@
Figure S3: Ablation study on influence of training-data diversity for denoising. Paired real data: The denoising network is trained using real data from the SID Sony training set [4], where some camera settings have very few unique scenes (e.g., the camera setting with ISO 6400 and ratio 300 has only one unique scenes). NoiseDiff: The denoising network is trained with our full set of synthetic data, where camera settings with limited unique scenes are augmented by scenes from other settings. Synthetic data is generated using the augmented scenes. NoiseDiff (Limited data): The denoising network is trained with limited synthetic data, where the number of scenes for each camera setting matches the original SID training set. The above three test samples from the SID testing set: (1) ISO 6400, ratio 300 (only one unique clean image in the original SID training set); (2) ISO 12800, ratio 100 (three unique clean images originally); and (3) ISO 800, ratio 100 (two unique clean images originally). It can be observed that the denoising performance improves with more synthetic data.
@@
11 More Discussions
11.1 Influence of Training-Data Diversity
@@
Standard Configuration. The original SID training set contains varying numbers of scenes across different camera settings, with some settings having very few scenes. For instance, the camera setting with ISO 6400 and ratio 300 has only one unique scenes. With our data generation model, however, we can leverage all the clean images in the SID training set to generate numerous training pairs for each camera setting. Specifically, if there are $K$ clean images in total, we can generate $K$ noisy images for each camera setting by conditioning our diffusion model on the camera setting embedding. Furthermore, since DDPM sampling is non-deterministic, different random seeds allow us to generate infinite noisy images from a single clean image. For efficiency, though, we generate a certain number of noisy images for each camera setting. For any setting with fewer than 6 unique clean images ($K_{i}<6$), we randomly sample $30-K_{i}$ additional clean images from other settings to augment the clean image number, then generate noisy images from both the original and augmented clean images. For settings with more than 6 clean images, we use only the original clean images for data generation.
Limited-Data Configuration. Here, we generate data without clean image augmentation. Specifically, for each camera setting with $K_{i}$ clean images in the SID training set, we generate noisy images using only these $K_{i}$ clean images. Therefore, the denoising network under this limited-data configuration uses the same number of scenes per camera setting as in the original SID training set.
@@
Fig. S3 presents a visual comparison using three test samples from the SID testing set: (1) ISO 6400, ratio 300 (only one unique clean image in the original SID training set); (2) ISO 12800, ratio 100 (three unique clean images originally); and (3) ISO 800, ratio 100 (two unique clean images originally). As shown, NoiseDiff produces better results than NoiseDiff (Limited data), benefiting from the additional synthetic training data. This further demonstrates that denoising performance improves with more synthetic data.
@@ 11.2 Sampling Method
11.3 Noise Decomposition
Implicit Noise Decomposition. While noise decomposition on real noise images is not possible due to the absence of ground-truth for the signal-dependent and signal-independent components, the conditional generation capability of our diffusion model allows us to perform implicit noise decomposition on the generated noise.
@@ 11.3 Noise Decomposition
We further analyze the statistics of different noise components. In Fig. S5 (g) and (j), we compare the variance and mean values of our generated noise with those of the real noise; they are calculated using the method described in Sec. 10. The results indicate that our method produces a noise distribution similar to the real noise. In Fig. S5 (h) and (k), we present the statistics of the signal-independent noise; they clearly show that the noise mean and variance are independent of the clean pixel values. Conversely, Fig. S5 (i) and (l) depict the statistics of the signal-dependent noise, demonstrating that both the noise variance and mean are correlated with the clean pixel values.
@@
Additional Evidence Supporting the Effectiveness of the Two-Branch Network. Meanwhile, we also apply the same decomposition process to the noise generated by the UNet-only architecture, with the results shown in Fig.S5 (m)-(u). It is evident that the signal-independent and signal-dependent components are not fully disentangled. For instance, in Fig.S5 (o), low-frequency noise is still present at the bottom of the signal-dependent component. This performance difference between the two architectures further highlights the effectiveness of the two-branch architecture in modeling signal-independent and signal-dependent noise, simplifying the network’s learning process.
TABLE S1: Ablation study on the dark shading correction (DSC) and shot noise augmentation (SNA) methods. We show the PSNR/SSIM results on the SID and ELD Sony test sets.
Model 1 Model 2 Model 3
Methods DSC
SNA
SID ×$\times 100$ 43.30 / 0.958 43.76 / 0.961 43.92 / 0.961
×$\times 250$ 40.53 / 0.944 41.26 / 0.945 41.28 / 0.946
×$\times 300$ 37.68 / 0.928 37.89 / 0.930 37.90 / 0.929
ELD ×$\times 100$ 45.79 / 0.972 46.51 / 0.979 46.95 / 0.978
×$\times 200$ 42.25 / 0.924 44.90 / 0.972 45.11 / 0.971
Figure S7: Histogram of noise values and Kullback-Leibler (KL) divergence between generated noise and real noise for four ablation variants: the UNet-only architecture, model without positional encoding (W/o PE), our full model with DDIM sampling (Full network (DDIM)) and our full model with DDPM sampling (Full network). Evaluations are conducted across three different ISO settings. The results produce by our full model are closest to the real noise distribution.
11.4 Noise Histogram
@@
11.5 Influenc of Dark Shading Correction and Shot Noise Augmentation
We employ dark shading correction (DSC) and shot noise augmentation (SNA) when training the denoising network of NoiseDiff* to ensure a fair comparison with LRD and PMN. As shown in Table S1, with the additional dark shading data, DSC significantly enhances denoising performance. SNA yields a smaller improvement. The combination of DSC and SNA achieves the best overall performance.
12 More Visual Comparisons
In this section, we provide more visual comparisons of denoising results from different networks trained on data synthesized using different methods. Fig. S8 - Fig. S11 are examples from the SID test set and Fig. S12 - Fig. S13 are examples from the ELD test set.
@@ 12 More Visual Comparisons
Our method excels at mitigating low-frequency noise, particularly visible in the third example of Fig. S8, the first examples of Fig. S9 and the second example of Fig. S10. Moreover, it enhances detail retention in denoising results, for example, producing clearer text compared to other methods, as shown in the third examples of Fig. S9, the first example of Fig. S10 and the first example of Fig. S13. Additionally, our method demonstrates effectiveness in handling row noise, as evidenced in the third examples of Fig. S12.
@@
Figure S8: Visual comparisons of denoising results from networks trained on data synthesized using different methods, alongside comparisons with a model trained on real clean-noisy pairs from the SID training set. All test samples are from the SID Sony test set. Zoom in for a clearer view.
@@
Figure S9: Visual comparisons of denoising results from networks trained on data synthesized using different methods, alongside comparisons with a model trained on real clean-noisy pairs from the SID training set. All test samples are from the SID Sony test set. Zoom in for a clearer view.
@@
Figure S10: Visual comparisons of denoising results from networks trained on data synthesized using different methods, alongside comparisons with a model trained on real clean-noisy pairs from the SID training set. All test samples are from the SID Sony test set. Zoom in for a clearer view.
@@
Figure S11: Visual comparisons of denoising results from networks trained on data synthesized using different methods, alongside comparisons with a model trained on real clean-noisy pairs from the SID training set. All test samples are from the SID Sony test set. Zoom in for a clearer view.
@@
Figure S12: Visual comparisons of denoising results from networks trained on data synthesized using different methods, alongside comparisons with a model trained on real clean-noisy pairs from the SID training set. All test samples are from the ELD test set. Zoom in for a clearer view.
@@
Figure S13: Visual comparisons of denoising results from networks trained on data synthesized using different methods, alongside comparisons with a model trained on real clean-noisy pairs from the SID training set. All test samples are from the ELD test set. Zoom in for a clearer view.